A sphere of radius ? is constrained to roll without slipping on the lower half of the inner surface of a hollow cylinder of inside radius R. Determine the Langrangian function, the equation of constraints, and Lagrange’s equation of motion. Find the frequency of small oscillations.

please explain what/why you're doing what you're doing. thank you
Answer #1

brion CR-18=10 Noslip condition CR ((RS) 0 7 sh-R/R-seso h = R(k 2 generalized coordinaty (0,0) but constraint reduces this tHence Lapo - Dom(R-2 52 mg ER-R-37 eso] de = m (R-8320 Therefore the equation of motion sont les como tmg CR-) smoco öt sgozo